OpenStudy (anonymous):

Solve for c. 2abc + d = 3

OpenStudy (apoorvk):

lol. subtract 'd' from both sides.

OpenStudy (apoorvk):

a,b,c,d are basically just terms. 2abc = 3 -d divide by '2ab' \[\frac{2abc}{2ab}=\frac{3-d}{2ab}\]

OpenStudy (apoorvk):

so what do you get now?

OpenStudy (apoorvk):

don't think what ab equals - the answer just be in terms of a,b,c, and d. that's it. \[\frac{\cancel{2ab}c}{\cancel{2ab}}=\frac{3-d}{2ab}\] this^ helps?

OpenStudy (anonymous):

c=idk

OpenStudy (apoorvk):

c = NOT 'idk' \[c=\frac{3-d}{2ab}\]

OpenStudy (anonymous):

ok thank you

OpenStudy (apoorvk):

that's it!^ don't get afraid of the 'alphabets' - they are just a form of denotion, used in place of the numbers.

OpenStudy (anonymous):

ok